# GEDAN: Learning the Edit Costs for Graph Edit Distance

### v1

@@ Appendix C Details of Edit Cost Optimization
For edit cost optimization we use 3 molecular datasets, varying the size and type of task required. The graphs are limited to 32 nodes, thus using the Gumbel-Sinkhorn 64x64 module, with the performance reported in Appendix E. We show the characteristics of the datasets used in Table 11.
@@
FreeSolv Mobley and Guthrie [2014] BBBP Martins et al. [2012] ZINC Gómez-Bombarelli et al. [2016] N. graphs Avg. nodes Avg. edges N. graphs Avg. nodes Avg. edges N. graphs Avg. nodes Avg. edges 642 8.72 16.78 1734 20.84 44.88 9615 22.78 48.91
FreeSolv Mobley and Guthrie [2014] is a small dataset that collects hydration free energies (HFE), both experimental and calculated, for small molecules in water. It is commonly used for the prediction of molecular properties in the field of computational chemistry. The choice of this dataset is due to its small size, especially in application contexts one does not have a lot of data, so we verified that the model would work in this context.
BBBP (Blood-Brain Barrier Penetration) Martins et al. [2012] is dataset that models the permeability of the blood-brain barrier, which can prevent the passage of the of drugs, hormones and neurotransmitters. The dataset tests the ability of a molecule to cross this barrier with binary labels. The dataset is not large to test the model in a real scenario of experimental data, where there are often not many.
ZINC Gómez-Bombarelli et al. [2016] is a large dataset of molecules, primarily used for virtual screening in drug discovery. The ZINC database contains millions of chemical compounds with information on their physicochemical properties, molecular structures, and commercial availability. We used a reduced version by testing the scalability of our model on a more complex feature.
C.1 Loss and Training Strategy
The training of the model S-GEDAN is based on the joint optimization of two loss functions:
@@ C.1 Loss and Training Strategy
C.2 Edit Cost Optimization Effect
In Fig. 567, we present the Principal Components Analysis (PCA) Dunteman [1989] obtained on the FreeSolv Mobley and Guthrie [2014], BBBP Martins et al. [2012] and ZINC Gómez-Bombarelli et al. [2016] datasets, respectively. PCA is performed on the 16 GED embedding values predicted by the networks repeating the process 3 different samplings (identify by letter (a), (b), (c) in the respectively figures). We present on the left the baseline version (U-GEDAN), in which the edit cost operations are uniform, on the right, the version optimized (S-GEDAN).
Each dot represents a graph while the color indicates the value of the associated label. As can be easily observed, PCA highlights the effect of optimization on the GED produced. The optimized methods show, in general, a better gradient transition, resulting in a more accurate alignment of target values.
Figure 5: PCA of the FreeSolv dataset Mobley and Guthrie [2014].
Figure 6: PCA of the BBBP dataset Martins et al. [2012].
Figure 7: PCA of the ZINC dataset Gómez-Bombarelli et al. [2016].
Appendix D Inference Time Analysis
@@
Figure 8: Time comparison in milliseconds of models inference on GED approximation dataset.
As shown in Fig.8, our model is significantly slower than the benchmarks. This slowdown is primarily due to an overhead introduced by matrix initialization rather than the number of trainable parameters.
Table 12 reports the average number of trainable parameters for each model. Our U-GEDAN model has fewer parameters than ERIC Zhuo and Tan [2022], while S-GEDAN has fewer parameters than GREED Ranjan et al. [2022]. Nevertheless, GEDAN models are slower in practice. Notably, models such as SimGNN Bai et al. [2019] and GraphEdX Jain et al. [2024] achieve excellent performance with a very small parameter count.
@@
SimGNN (2019) GREED (2022) ERIC (2022) GraphEdX (2024) U-GEDAN S-GEDAN  Bai et al. [2019]  Ranjan et al. [2022]  Zhuo and Tan [2022]  Jain et al. [2024] (Our) (Our) 9249 151553 29157 5785 15306 53182
@@
Inference time for molecular prediction is also compared against GraphMaskExplainer Schlichtkrull et al. [2021]. Both GEDAN and the reference GNN have 4 layers to ensure that prediction time is not biased by depth. In the direct comparison, first column of Fig.2 (a), only GEDAN’s result is reported, as GraphMaskExplainer cannot be applied in this case. The other two columns refer to predictions shown in Fig.2 (b) and (c), where GEDAN is faster.
Figure 10: Time comparison in seconds of prediction of GEDAN molecules compared with GNN analyzed by GraphMaskExplainer Schlichtkrull et al. [2021].
Appendix E Model Details
E.1 Gumbel-Sinkhorn
To obtain a good $\mathbf{P}$, we modify the version of the original Gumbel-Sinkhorn network Mena et al. [2018], Shah et al. [2024], using the Truncated Gumbel Neamah and Qasim [2021] to reduce the gradient variance and add stability. We then sample with the Gumbel-Max Trick Huijben et al. [2023] while maintaining differentiability. Finally we stabilize the normalization of Sinkhorn Sinkhorn [1964] with an early stop, improving efficiency and limiting gradient vanishing LeCun et al. [1998].
The Gumbel-Sinkhorn network has been trained for 200 epochs on matrices of sizes $32\times 32$, $64\times 64$ and $128\times 128$. The matrices used for training are randomly generated, with 250000 samples for each of the three dimensions. For each matrix, Linear Sum Assignment (LSA) is calculated and used as a reference label. Optimization is performed by minimizing the Huber Loss, with a learning rate of 0.001 and using the Adam optimizer. Training also includes a curriculum learning phase, during which many matrices with zero values on the diagonal are initially proposed, a strategy that facilitated model convergence. In addition, an early stopping mechanism is applied on the Sinkhorn process, with a limit of 50 iterations, where the stopping condition is determined by a difference between consecutive iterations of less than 1e-4.
To test the performance of the Gumbel-Sinkhorn network, 10000 new random matrices are generated, for which the LSA varies between 0.001 and 199.750, with an average of 77.375. The overall performance of the model calculated on the $64\times 64$ matrices using high temperatures has a $\tau$ value of 0.991, a $R^{2}$ coefficient of 0.993 and a Root Mean Square Error (RMSE) of 3.940. These configurations are kept stable in the pre-trained versions in the U-GEDAN and S-GEDAN models.
E.2 U-GEDAN and S-GEDAN
As for the functions $f^{(k)}$, we concatenate the representations of the two nodes involved and provide the result as input to an MLP. This MLP employs a ReLU as an intermediate activation function and applies a softplus function with parameter $\beta=5$ on the final output, as described in the main body of the text.
For U-GEDAN and S-GEDAN training of the GED approximation we use a NVIDIA A100. Cost optimization training is performed on a cluster using two NVIDIA H100s. The model implementation is carried out using PyTorch (version 2.2.0) and the PyTorch Geometric library (version 2.4.0). Hardware acceleration is enabled via CUDA (version 12.1). The source code for the model is publicly available at https://github.com/GEDAN-neurips2025/GEDAN.
Appendix F Predicted Molecules
@@ F.1 Direct Comparison
Figure 11: Direct comparison between pairs of similar molecules. Each panel (a–c) shows two molecules differing by a small chemical modification.
F.2 Symmetric Molecules
Molecular symmetry poses a unique challenge for predictive models because it requires invariance or equivariance with respect to permutations of atoms or substructures. In this section, we evaluate the model’s ability to produce predictions that respect molecular symmetry, both at the level of global structure and local atoms. The examples in Fig. 12 demonstrate that our model maintains more symmetric predictions than GraphMaskExplainer Schlichtkrull et al. [2021].
Figure 12: Model predictions on symmetric molecules. Each panel (a – e) illustrates the model’s ability to produce equivalent results for symmetrically equivalent atoms or substructures.
F.3 Chemically Consistent Molecules
To assess the model’s ability to generalize to structurally different but chemically similar compounds, we evaluate its predictions on a set of molecules that share some functional properties. This analysis complements the main examples by demonstrating that the model preserves chemically meaningful behavior beyond simple or symmetric cases. As shown in Fig. 13, the model produces consistent results.

### v2

@@ Appendix E Details of Edit Cost Optimization
For edit cost optimization we use two molecular datasets, varying the size and type of task required. The graphs are limited to 32 nodes, thus using the Gumbel-Sinkhorn $64\times 64$ module, with the performance reported in Appendix C. We show the characteristics of the datasets used in Table 15.
@@
FreeSolv (Mobley & Guthrie, 2014) BBBP (Martins et al., 2012) N. graphs Avg. nodes Avg. edges N. graphs Avg. nodes Avg. edges 642 8.72 16.78 1734 20.84 44.88
FreeSolv (Mobley & Guthrie, 2014) is a small dataset that collects hydration free energies (HFE), both experimental and calculated, for small molecules in water. It is commonly used for the prediction of molecular properties in the field of computational chemistry. The choice of this dataset is due to its small size, especially in application contexts one does not have a lot of data, so we verified that the model would work in this context.
Blood-Brain Barrier Penetration (BBBP) (Martins et al., 2012) is dataset that models the permeability of the blood-brain barrier, which can prevent the passage of the of drugs, hormones and neurotransmitters. The dataset tests the ability of a molecule to cross this barrier with binary labels. The dataset is not large to test the model in a real scenario of experimental data, where there are often not many.
E.1 Loss and Training Strategy
The training of GEDAN, aimed at aligning the GED with the functional distance and thus learning the optimal set of edit costs, relies on the joint optimization of two loss functions:
@@ E.1 Loss and Training Strategy
E.2 Edit Cost Optimization Effect
In Figs.16 and17, we present the Principal Component Analysis (PCA)(Dunteman, 1989) results for the FreeSolv and BBBP datasets, respectively. PCA is applied to the 16 GED embedding values predicted by the networks GEDAN($\lambda=1$), GMSM ($\epsilon\geq 0$(Pellizzoni et al., 2024), GMSM ($\epsilon=0$(Pellizzoni et al., 2024), and GEDAN ($\lambda=0$). All models employ 5 layers to ensure that both architectures perceive the same graph depth. For GMSM (Pellizzoni et al., 2024), we use the retrieve method, which provides a similarity value used to construct the embeddings, whereas GEDAN directly outputs the GED, which we employ for PCA. For GEDAN ($\lambda=1$), we adopt the uniform fixed cost variant with costs equal to 1.
In Figs.16 and17, each point corresponds to a graph, and its color indicates the associated label value. PCA clearly illustrates the effect of optimization on the predicted GED: optimized methods generally exhibit smoother gradient transitions, leading to a more accurate alignment with the target values. This additional experiment further demonstrates that the use of uniform costs is often poorly representative for many real-world scenarios.
Figure 16: PCA of the FreeSolv dataset (Mobley & Guthrie, 2014).
Figure 17: PCA of the BBBP dataset (Martins et al., 2012).
Appendix F Inference Time Analysis
@@
Figure 18: Time comparison in milliseconds on a logarithmic scale of model inference on the GED approximation dataset.
As shown in Fig. 18, our model is considerably slower than supervised models. This slowdown is primarily due to the overhead introduced by matrix initialization rather than the number of trainable parameters. In contrast, EUGENE (Bommakanti et al., 2024), which is also unsupervised, is significantly slower, taking several seconds to produce a single output. Its CPU-based optimization does not allow for acceleration of the code. Moreover, as illustrated in Figs. 12 and 13, we can train GEDAN for approximately 20 epochs before EUGENE (Bommakanti et al., 2024) completes prediction on the same dataset.
Table 16 reports the average number of trainable parameters for each model. Although GEDAN has fewer parameters than ERIC (Zhuo & Tan, 2022), EGSC (Qin et al., 2021), SimGNN (Bai et al., 2019a), and GREED (Ranjan et al., 2022), it is slower in practice. Notably, GraphEdX (Jain et al., 2024) achieves strong performance with few parameters. EUGENE (Bommakanti et al., 2024), being non-neural, has no trainable parameters. For all other models, we tune hyperparameters to maximize performance on the dataset, taking into account the differences in parameter counts.
@@
SimGNN EGSC GREED ERIC GraphEdX GEDAN  (Bai et al., 2019a)  (Qin et al., 2021)  (Ranjan et al., 2022)  (Zhuo & Tan, 2022)  (Jain et al., 2024) ($\lambda=1$) ($\lambda=0$) 34,894 107,316 151,553 59,156 28,535 15,306 53,182
@@
The inference time for molecular interpretation is compared with other models on a logarithmic scale. In Fig. 3 (a), only GEDAN and GMSM (Pellizzoni et al., 2024) can perform the task, achieving comparable times, with initialization overhead being the main cost. Figures 3 (b) and (c) show inference times for all models. GraphMaskExplainer (Schlichtkrull et al., 2021) and PGExplainer (Luo et al., 2020) include 250 training epochs in their reported times. GATv2 (Brody et al., 2022) is the fastest, as inference only requires visualizing attention weights. Overall, GEDAN is slower than GMSM (Pellizzoni et al., 2024) but faster than GraphMaskExplainer (Schlichtkrull et al., 2021) and PGExplainer (Luo et al., 2020), due to the handling of the maximum cost assigned to node types, which improves local precision at a computational cost.
Figure 20: Inference time for molecular interpretation across different models. The direct comparison corresponds to Fig. 3 (a), the first molecule to Fig. 3 (b), and the second molecule to Fig. 3 (c).
Appendix G Predicted Molecules
@@ G.1 Direct Comparison
Figure 21: Direct comparison between pairs of similar molecules. Each panel (a – d) shows two molecules differing by a small chemical modification.
G.2 Symmetric Molecules
Molecular symmetry poses a unique challenge for predictive models because it requires invariance or equivariance with respect to permutations of atoms or substructures. The comparisons in which a single molecule is compared to all others are done one by one. After that, we calculate the average costs at the node level, which gives us a measure of the average transformation cost. This process allows us to determine, on average, which atoms are most important. In Fig. 22, we evaluate the models’ ability to produce predictions that respect molecular symmetry, both at the level of global structure and local atoms.
Figure 22: Model predictions on symmetric molecules. Each panel (a – d) illustrates the model’s ability to produce equivalent results for symmetrically equivalent atoms or substructures.
G.3 Chemically Consistent Molecules
To assess the model’s ability to generalize to structurally different but chemically similar compounds, we evaluate its predictions on a set of molecules that share some functional properties. In Fig. 23, we complement the main examples by demonstrating that the model preserves chemically meaningful behavior beyond simple or symmetric cases.
@@ G.3 Chemically Consistent Molecules
Appendix H Specifications on the Limitation of the Number of Nodes
A major limitation of our method concerns the maximum size that can be handled by the Gumbel-Sinkhorn matrices, which introduce quadratic computational complexity. This aspect limits the model’s ability to generalize to large-sized graphs. For this reason, we restricted the application of our approach to small molecules, excluding more complex structures such as peptides or proteins. In addition to the larger size of the graphs, the nature of the problem also changes in these cases, since macro-molecules are strongly affected by three-dimensional folding. Next, we present the maximum size of the graphs that our model can handle and the coverage with respect to the main benchmark datasets currently available online.
Table 17: Size distribution of molecular graphs present in the principal public benchmarks
Nodes Maximum size of GEDAN
Datasets Total Avg Std Max 64 128 256
AQSOL (Dwivedi et al., 2023a) 7,836 15.42 8.63 156 7,812 7,834 7,836
MD17 (Christensen & von Lilienfeld, 2020) 576,583 576,583 576,583 576,583
MoleculeNet (Wu et al., 2017) 596,815 595,116 596,659 596,764
PCQM4Mv2 (Hu et al., 2021) 3,378,606 14.12 2.47 20 3,378,606 3,378,606 3,378,606
QM7b (Wu et al., 2017) 7211 15.42 2.69 23 7,211 7,211 7,211
QM9 (Wu et al., 2017) 130,831 18.03 2.94 29 130,831 130,831 130,831
TUDataset (Morris et al., 2020) 4,789,598 4,723,791 4,785,518 4,789,594
ZINC (Gómez-Bombarelli et al., 2016) 220,011 23.15 4.51 38 220,011 220,011 220,011
Total 9,707,491 9,639,961 9,703,253 9,707,436
Total (%) 99.30% 99.96% 100.00%